1.1 find parametric equations of the line that passes through the point P =(2, 0,-1) and is parallel to the vector n =<2, 1, 3>

1.2 find paramedic equations of the line that passes through the points A= (1, 2, - 3) and B =(7, 2, - 4).

1.3 find paramedic equations for the line of intersection of the planes - 5x + y - 2z =3 and 2x - 3y + 5z =-7



1
Expert's answer
2021-07-16T08:01:30-0400

Answer:-

1.1)


P = (2, 0,-1)


vector n =<2, 1, 3>


Let the parameter of the line be t . Since the line is parallel to vector n, so the direction ratios of the line to be found will also be parallel to the vector n


The equation of the line is given by


"\dfrac{x - x_1}{a} = \dfrac{y-y_1}{b} = \dfrac{z-z_1}{c} = t"



So, on substituting the values in the above equation, we have


"\dfrac{x - 2}{2} = \dfrac{y-0}{1} = \dfrac{z+1}{3} = t"



From the above equation we get the parametric equation of line as


x = 2t + 2 , y = t, z = 3t - 1





1.2)



A= (1, 2, - 3) and B =(7, 2, - 4).


The equation of the line is given by


"\dfrac{x - x_1}{a} = \dfrac{y-y_1}{b} = \dfrac{z-z_1}{c} = t"



Here,


a = x2 - x= 7 - 1 = 6

b = y2 - y= 2 - 2 = 0

c = z2 - z= -4 + 3 = -1



Hence, on substituting the values we have


"\dfrac{x - 1}{6} = \dfrac{y-2}{0} = \dfrac{z+3}{-1} = t"


We get the parametric equation of line as


x = 6t + 1, y = 2, z = - t - 3





1.3)


- 5x + y - 2z = 3

and

2x - 3y + 5z = - 7


To find the parametric equation of line we will solve the given system of planes.


Let z = t, where t is our free parameter.


So, the equation of the plane takes the form


-5x + y = 3 + 2t

2x - 3y = -7 - 5t


On solving the above equations we get


x = "\dfrac{-1}{13}(2+t)"


y = "\dfrac{1}{13}(29 + 2t)"


So, the parametric equation of the line of intersection of the given planes is



x = "\dfrac{-1}{13}(2+t)", y = "\dfrac{1}{13}(29 + 2t)" , z = t